OpenStudy (anonymous):

Identify the zeros of the function y-(x-8)(x-15)

OpenStudy (anonymous):

If it's, y=(x-8)(x-15) x = 8 and x = 15

OpenStudy (anonymous):

Set y = to zero, and solve for x. 0=(x-8)(x-15) x-8 = 0 x-15 = 0